125ml of 1.00M HCl is added to 100ml of 1.00M NaOH. What is the pH of the resulting solution? (Please show your work.)
1
Expert's answer
2020-10-16T08:55:32-0400

Millimoles of "HCl= 125\times 1.00= 125"

Millimoles of "NaOH= 100\times 1.00= 100" Sodium hydroxide reacts with hydrochloric acid in equal amount. So, "100" millimoles will be neutralised of both reactants.

So , "25" millimoles of "HCl" will remain unreacted .

So moles of "H^+" in solution will be "25\times 10^{-3}= 0.025"

Total volume of solution "= 125+100 \ ml = 225 \ ml"

Concentration of "H^+ = \frac{0.025}{225\times 10^{-3}}=0.11\ M"

"pH" of solutiob"=-log[H^+] = -log(0.11) = 0.95"